Can you please help me solve the proportion 8/x+3= 1/x +1 (8:x+3=1:x=1
This is what I have so far: I got x=1 by isolating x, but my paper says to factor the equation

Answer by stanbon(75887)

8/x+3= 1/(x +1)
-------
Cross-multiply to get:
8(x+1) = (x+3)
---
8x + 8 = x+3
7x = -5
x = -5/7
---------
Cheers,
Stan H.
=============